Languages, Inc. manufactures hand-held computers that translate between two languages. Based on their market research, they have developed computers for French/English, German/English, and Spanish/English. The process for making the computers will have fixed costs of $2,000,000 per year and variable costs of $50 per computer. The company believes that it can sell at least 40,000 computers per year.

a) What should the price per computer be if the company wants to break even at a volume of 40,000 computers per year?

b) If they sell 60,000 computers at a price of $90 per computer, what will the contribution to profit be?

please write correct and clear answer
Answer #1

a) Let the price of each computer at a volume of 40000 be X, then:

2000000 + 50*40000 = 40000*X

40000*X = 4000000

X = 100

In order to break even at a volume of 40,000 the price per computer needs to be $100

b) if the volume = 60000

Selling Price = $90

Total expense

= 2000000 + 60000*50

= 5000000

Total Revenue

= 60000*90

= $5400000

Profit = 5400000-5000000

= $400000